How many three digit numbers can be formed from the set 0-9 if repetition are allowed and the number must be
a) even
b) divisible by 5?
c)divisible by 10?

Answer by Edwin McCravy(20054)

How many three digit numbers can be formed from the set 0-9 if repetition are allowed and the number must be
a) even
Choose the first digit as any of these 9 ways: {1,2,3,4,5,6,7,8,9}
Choose the second digit as any of these 10 ways: {0,1,2,3,4,5,6,7,8,9}
Choose the third digit as any of these 5 ways: {0,2,4,6,8}

Answer: 9×10×5 = 450 ways.

b) divisible by 5?
Choose the first digit as any of these 9 ways: {1,2,3,4,5,6,7,8,9}
Choose the second digit as any of these 10 ways: {0,1,2,3,4,5,6,7,8,9}
Choose the third digit as either of these 2 ways: {0,5}

Answer: 9×10×2 = 180 ways.

c)divisible by 10?
Choose the first digit as any of these 9 ways: {1,2,3,4,5,6,7,8,9}
Choose the second digit as any of these 10 ways: {0,1,2,3,4,5,6,7,8,9}
Choose the third digit as this 1 way: {0}

Answer: 9×10×1 = 90 ways.

Edwin